# Enhancing Social Decision-Making of Autonomous Vehicles: A Mixed-Strategy Game Approach With Interaction Orientation Identification

### v1

@@ I Introduction
Current autonomous driving decision-making algorithms include rule-based [7], deep learning-based [8], and game theory-based methods [3, 9]. Significant research has been conducted on enhancing the social intent understanding and response capabilities of autonomous driving using these methods [10, 11, 12].
@@ III-B Framework Overview
The social decision-making game framework we propose is outlined in Fig. 2. It is composed of three integral modules: Interaction Orientation Identification, Mixed-Strategy Game Modeling, and Expert Mode Learning. In the Interaction Orientation Identification phase, the framework quantifies the decision-making tendencies of the interacting subjects utilizing the concept of Interaction Orientation. Following this, a Mixed-Strategy Game Model is developed. This model incorporates the evolution of future states and integrates a utility function that balances safety, operational efficiency, and the unpredictability of environmental factors.
To refine the social decision-making capabilities, a dynamic optimization process is employed, which is designed to learn and adapt interaction patterns based on real-world data. This process involves building models that learn from human driving experts and storing these insights within an expert strategy library. In the application phase, the framework utilizes the interaction orientation identification module to discern the inclinations of interacting subjects and accesses the corresponding expert strategies. This facilitates the efficient resolution of the mixed-strategy social game model in real-world scenarios.
@@ IV-A1 Modeling for Interaction Environment State
Subsequently, we further use the Softmax function [205] to assign weights to $\Delta TTCP_{norm}$ and $a_{c_{norm}}$, expressing them as a Interaction Transient State Index (ITSI):
@@ IV-A2 Modeling for Trajectories Motion Feature
If both the straight vehicle and the left-turning vehicle travel along the centerline of their virtual lanes, with the bottom-left coordinate of the occupied area as $t_{1},S_{1}$ and the top-right coordinate as $t_{2},S-2$, the coordinates are determined as follows:
@@ IV-B2 Payoff Function Design
where $TTCP(i,0)$ is the time required for the participant to pass through the conflict point under the current state, and $TTCP_{i}^{S}$ is the time required under decision $S$. The calculation method is as follows:
@@ IV-C Expert Mode Learning
1   Pre-classify the expert data set $\mathcal{D}$ based on interaction tendencies to obtain $\mathcal{D}=\{d_{1},d_{2},d_{3}\}$;
@@ V Data-based Experiments and Analysis
V-A Dataset
For model learning and validation, the Argoverse2 Motion Dataset [41] and SinD datasets [42] are employed. The Argoverse2 Motion Dataset, amassed by Argo AI across six cities in the United States, encompasses approximately 763 hours and includes over 440,000 scene groups rich in interactive scenarios. The SinD dataset, gathered by Tsinghua University at an unsignalized intersection in Tianjin, China, documents the trajectory data of more than 4,800 vehicles. From the Argoverse2 dataset, we select strong interaction scenarios involving left-turn and straight-on traffic at unsignalized intersections to validate the interaction orientation identification module. In the SinD dataset, 268 interaction events are extracted, which include 132 scenarios where left-turning vehicles take precedence, 136 where they yield, and the number of segments for straight-on traffic yielding, exhibiting no clear tendency, and proceeding first are 421, 800, and 362, respectively. This data serves for expert strategy learning.
@@ VII Conclusion
In the future, we aim to further extend our method to more complex and dynamic scenarios, such as considering the social decision-making of AVs when facing multiple interaction partners or vehicle groups, enhancing the adaptability and generalizability of our algorithm.

### v2

@@ I Introduction
Current autonomous driving decision-making algorithms include rule-based [7], deep learning-based [8, 9], and game theory-based methods [3, 10]. Significant research has been conducted on enhancing the social intent understanding and response capabilities of autonomous driving using these methods [11, 12, 13].
@@ III-B Framework Overview
Our proposed social decision-making game framework primarily consists of three modules: driver social preference identification, mixed-strategy game modeling, and expert mode learning. In the driver social preference identification module, we quantify the social decision-making tendencies of interaction objects through interaction tendencies. Subsequently, we construct a mixed-strategy game module that considers future state evolution and design a benefit function that accounts for safety, efficiency, and environmental random disturbances, aiming at achieving safe and efficient interaction decisions.
To enhance the generalization ability and adaptability of social decision-making, we design an expert mode learning module. This module aims to learn expert modes based on real-world data and dynamically adjusts interaction modes during the decision-making process. We develop an expert strategy database where the framework constructs an optimized model to learn human expert strategies and saves the learning outcomes to the expert strategy library. During the application process of the decision model, we utilize the social tendency module to identify the tendencies of interaction objects and search for and load the most stylistically similar expert mode parameters from the expert strategy library, enabling efficient solving of the mixed-strategy social game model.
@@ IV-A1 Modeling for Interaction Environment State
Subsequently, we further use the Softmax function to assign weights to $\Delta TTCP_{norm}$ and $a_{c_{norm}}$, expressing them as a Interaction Transient State Index (ITSI):
@@ IV-A2 Modeling for Trajectories Motion Feature
As illustrated in Fig.4(a), we have depicted the ST trajectory graph with the straight-going vehicle as the subject. The occupied area represents the temporal and spatial relationship of the straight-going vehicle appearing on its ST trajectory. The red area in the graph indicates the occupied area by the left-turning vehicle, which is determined by the size and instantaneous speed of the left-turning vehicle. As shown in Fig.4(a), the ST diagram is pictured with the main view of straight-going vehicle. The red area denotes the occupied area of left-turn vehicle, which is determined by the size and constant speed of the left-turn vehicle. If both the straight vehicle and the left-turning vehicle travel along the centerline of their virtual lanes, with the bottom-left coordinate of the occupied area as $t_{1},S_{1}$ and the top-right coordinate as $t_{2},S_{2}$, the coordinates are determined as follows:
@@
where the minimum longitudinal distance $S_{max}$ and the maximum longitudinal distance $S_{min}$ are determined based on Equation (15), with $\Delta S_{long}$ being the actual longitudinal displacement distance produced by the real trajectory within the time domain $[t_{0},t_{s}]$.
@@ IV-B2 Payoff Function Design
Here, $TTCP_{(i,0)}$ is the time required for participant $i$ to pass the conflict point under the current state, and $TTCP_{i}^{S}$ is the time required under decision $S$. The calculation of $TTCP_{i}^{S}$ needs to consider the impact of different decisions on behavior: if participant $i$ chooses to preempt, they move to the conflict point with uniform acceleration $a_{i}^{preempt}$; if yielding, they decelerate uniformly to a stop with $a_{i}^{yield}$, wait for the interaction partner $-i$ to pass, and then proceed with $a_{i}^{preempt}$. Thus, $TTCP_{i}^{S}$ is represented as:
@@ IV-C Expert Mode Learning
1
2Pre-classify the expert data set $\mathcal{D}$ based on interaction tendencies to obtain $\mathcal{D}=\{d_{1},d_{2},d_{3}\}$;
@@ V Data-based Experiments and Analysis
V-A Dataset and Parameter Setting
Dataset. For model learning and validation, the Argoverse2 Motion Dataset [43] and SinD datasets [44] are employed. The Argoverse2 Motion Dataset, amassed by Argo AI across six cities in the United States, encompasses approximately 763 hours and includes over 440,000 scene groups rich in interactive scenarios. The SinD dataset, gathered by Tsinghua University at an unsignalized intersection in Tianjin, China, documents the trajectory data of more than 4,800 vehicles. From the Argoverse2 dataset, we select strong interaction scenarios involving left-turn and straight-on traffic at unsignalized intersections to validate the interaction orientation identification module. In the SinD dataset, 268 interaction events are extracted, which include 132 scenarios where left-turning vehicles take precedence, 136 where they yield, and the number of segments for straight-on traffic yielding, exhibiting no clear tendency, and proceeding first are 421, 800, and 362, respectively. This data serves for expert strategy learning.
[…]
@@ VII Conclusion
In future endeavors, our objective is to broaden the application of our methodology to encompass more intricate and dynamic contexts. This expansion includes addressing the social decision-making processes of AVs when confronted with multiple interacting entities or conglomerations of vehicles. Such an advancement aims to augment the adaptability and generalizability of our algorithm significantly. Furthermore, we intend to undertake the extraction and assimilation of prior knowledge and potential norms governing social interactions from an extensive compilation of datasets derived from real-world experts. This initiative is designed to enrich and enhance the capabilities of autonomous driving algorithms in social decision-making contexts.